OpenStudy (anonymous):

Factor the polynomial completely. (If the expression is not factorable using integers, enter PRIME.) 6x^6 − 6 my answer (2x^3-2)(3x^3+4x^3+3) what did I do wrong?

OpenStudy (anonymous):

6 (x^6 -1) x^6 - 1 = (-1+x) (1+x) (1-x+x^2) (1+x+x^2) 6x^6 - 6 = 6(-1+x) (1+x) (1-x+x^2) (1+x+x^2)

OpenStudy (anonymous):

\[6x^6 - 6 = 6(-1+x) (1+x) (1-x+x^2) (1+x+x^2)\]

jhonyy9 (jhonyy9):

\[6x ^{6}-6 =6(x ^{6}-1)=6(x ^{2})^{3}-1=6(x-1)(x ^{2}+x+1)\]

jhonyy9 (jhonyy9):

sorry

jhonyy9 (jhonyy9):

\[6(x ^{2}-1)(x ^{4}+x ^{2}+1)\]

jhonyy9 (jhonyy9):

this will be right , correct

jhonyy9 (jhonyy9):

the last answer

jhonyy9 (jhonyy9):

pa

jhonyy9 (jhonyy9):

pard

jhonyy9 (jhonyy9):

pardon

jhonyy9 (jhonyy9):

hope so much that is understandably

OpenStudy (anonymous):

(x^2−1) = (x+1)(x-1) (1-x+x^2) (1+x+x^2) = (x^4 + x^2 +1)

OpenStudy (anonymous):

It's the same as my answer, except my answer is more simplified

jhonyy9 (jhonyy9):

no moneybird us formula a3 -b3 =(a-b)(a2+ab+b2)

jhonyy9 (jhonyy9):

but you can using again the same 6(x6-1)=6((x3)2 -1 =6(x3-1)(x3+1) what you can continue

jhonyy9 (jhonyy9):

and finaly will get what i have wrote firstly

OpenStudy (anonymous):

No there's a formula for all n numbers (a^n - b^n = (a-b)(a^(n-1) + a^(n-2)b...b^(n-1)

OpenStudy (anonymous):

most regular schools won't teach you that. It doesn't mean this property does not exist

jhonyy9 (jhonyy9):

the right answer is 6(x-1)(x+1)(x4+x2+1)

OpenStudy (anonymous):

no. You can still simplify x^4+x^2+1

jhonyy9 (jhonyy9):

how ?

OpenStudy (anonymous):

ok let x^2 = a x^4+x^2+1 = a^2 + a + 1 factor this and plug x^2 back after

jhonyy9 (jhonyy9):

h

jhonyy9 (jhonyy9):

how ?

jhonyy9 (jhonyy9):

can you make it ?

jhonyy9 (jhonyy9):

camon please i like seeing how

jhonyy9 (jhonyy9):

how you can factorizing a2+a+1

OpenStudy (anonymous):

(1-x+x^2) (1+x+x^2) if you expand this, you will get x^4+x^2+1)

jhonyy9 (jhonyy9):

ok bye

jhonyy9 (jhonyy9):

thank you